AnimeGhoul8863:

The table below shows the squares of different numbers: Number (x) 2 −2 3 −3 Square of the number (y) 4 4 9 9 Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a rowboat for x hours is shown below: f(x) = 20 + 3x What is the value of f(150), and what does f(150) represent? (5 points) -----

AnimeGhoul8863:

Part A: Yes The table represents y as a function of x Because the function is f(x) = x^2 X= inputs Y= Outputs For example f(x) = x^2 f(-3) = (-3)^2 input= x = -3 f(-3) = 9 output = 9 Part B: 150=x f(150) = 20+3*150 f(150) = 20+300 f(150) = 320 it represents 150 hours of renting a row boat = 320 dollars

AnimeGhoul8863:

the answer @dude

dude:

Looks right!

AnimeGhoul8863:

i really have to get like 100 on this or ill fail please make sure i do good on these

AnimeGhoul8863:

this one i havent answered yet tho so can you help me with it

dude:

Well for part 1 \[x^2\\ ~(2)^2=4\\ ~(-2)^2=4\\ Likewise~with~3~and~-3\]

AnimeGhoul8863:

what

dude:

Part A is right lol I just rewrote your work

AnimeGhoul8863:

this is the question

AnimeGhoul8863:

Sam is observing the velocity of a car at different times. After three hours, the velocity of the car is 51 km/h. After five hours, the velocity of the car is 59 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first six hours? (5 points)

dude:

Well what are the x and y values?

AnimeGhoul8863:

x=51 y=59?

dude:

I meant what do they represent?

AnimeGhoul8863:

ummmmmmmmmmm

AnimeGhoul8863:

my mind is literally blank i dont know

dude:

What are we measuring?

AnimeGhoul8863:

km

dude:

Yes speed and?

AnimeGhoul8863:

distance

dude:

No \[Hint:~"After~{three}~\underline{\color{red}{hours}},~the~velocity~of~the~car~is~51~km/h."\]

AnimeGhoul8863:

velocity

dude:

Hours is not velocity x'D

AnimeGhoul8863:

oh i didnt see the under line

AnimeGhoul8863:

so speed and hours

dude:

Yes, "hours" was underlined and colored i don't understand how you didn't see it x'D

AnimeGhoul8863:

look man ive been doing test since 9 Am give me a little slack im exhausted

dude:

~Hours is the dependent variable and speed is the dependent

AnimeGhoul8863:

so the variables would be S and H

AnimeGhoul8863:

am i right

dude:

Variable letters don't really matter but sure

AnimeGhoul8863:

the question ask for Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times.

AnimeGhoul8863:

ok lets focus its 7 pm and i want to be done with school

dude:

Lets find the distance between these points

AnimeGhoul8863:

ok good

dude:

\[\frac{y_2-y_1}{x_2-x_1}=\frac{59-51}{5-3}=?\]

AnimeGhoul8863:

=4

dude:

Right, that was the slope formula, so \[y=mx+b\\ y=4x+b\] To find b, we can substitute a point given in the question

dude:

\[y=4x+b~~~~"After~{three}~{{hours}},~the~velocity~of~the~car~is~51~km/h.""\\ (x,y) ~x=hours,~~~y=velocity\\ (3,51)\\ \color{green}{51=4(3)+b}\\ 51=12+b\\ Subtract~12~from~both~sides\\ 39=b\\~\\ So~\color{red}{y=4x+39}\]

AnimeGhoul8863:

soooo the answer for part A is x=hours down to +39

dude:

No, the equation is in red!! ;-;

AnimeGhoul8863:

ik but you have to show work

dude:

I just did!!!! e.e

AnimeGhoul8863:

you said "the equation is in red" like that was the only answer i said i neede dto add work so i am

AnimeGhoul8863:

any way i need part b and one more question and im done

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude \[y=4x+b~~~~"After~{three}~{{hours}},~the~velocity~of~the~car~is~51~km/h.""\\ (x,y) ~x=hours,~~~y=velocity\\ (3,51)\\ \color{green}{51=4(3)+b}\\ 51=12+b\\ Subtract~12~from~both~sides\\ 39=b\\~\\ So~\color{red}{y=4x+39}\] \(\color{#0cbb34}{\text{End of Quote}}\) The work is right above that line e.e

AnimeGhoul8863:

so its 6,63

dude:

Do you understand what it means?

AnimeGhoul8863:

yes

AnimeGhoul8863:

last question

AnimeGhoul8863:

The coordinate plane below represents a town. Points A through F are farms in the town. graph of coordinate plane. Point A is at 2, negative 3. Point B is at negative 3, negative 4. Point C is at negative 4, 2. Point D is at 2, 4. Point E is at 3, 1. Point F is at negative 2, 3. Part A: Using the graph above, create a system of inequalities that only contains points C and F in the overlapping shaded regions. Explain how the lines will be graphed and shaded on the coordinate grid above. (5 points) Part B: Explain how to verify that the points C and F are solutions to the system of inequalities created in Part A. (3 points) Part C: Chickens can only be raised in the area defined by y < 5x − 3. Explain how you can identify farms in which chickens can be raised. (2 points)